x/3- the square root of 14=5/2 what is x

Answer by stanbon(75887)

x/3- the square root of 14=5/2 what is x
----
(x/3) - sqrt(14) = (5/2)
Multiply thru by 6 to get:
2x - 6sqrt(14) = 15
2x = 6sqrt(14) - 15
x = (1/2)(6sqrt(14) - 15)
==============================
Cheers,
Stan H.